A sample of sodium azide (NaN 3), a compound used in automobile air bags was thermally decomposed and 15.3 mL of nitrogen gas was collected at 25°C and 755 torr. How many grams of nitrogen were collected?
Answer #1

For an ideal gas,

PV=nRT P=Pressure=(755/760) atm

= 0.9934 atm

V= Volume =15.3 ml = (15.3/1000) lit

= 0.0153 lit

R= ideal gas constant = 0.082 lit-atm mol-1​​-1k-1

T= Temperature = 25°C = (25+273) = 298 k

Now, for an ideal gas equation,

n = PV/RT

= (0.9934×0.0153)/(0.082×298)

= 0.000622 mol

Now, molar mass of Nitrogen = 28 g/ mol

So, 1 mol N​​​​​​2 contains 28 g of N2

0.000622 mol N2 contain (0.000622 × 28) g

= 0.0174 g

Therefore, 0.0174 g Nitrogen were collected.